What is the concentration in percentage by mass, of each of the following solutions?

a.      3.0 g of C2H5OH with enough H2O to make 400.00 g of solution

b.      40 g of KCl plus 180 g H2O 


1
Expert's answer
2021-02-11T04:46:53-0500

a.    

Total mass of the solution = 400 g

Mass of C2H5OH = 3.0 g

Mass of water = (400–3)g

              = 397 g


So, 

397 g water contain 3 g of C2H5OH

1 g water contain = (3/397)g

 100 g water contain = (3×100/397) g

                   = 0.756 g

                   

Hence, concentration in mass percentage of C2H5OH = 0.756 %

   


b.  

  40 g of KCl plus 180 g H2O

180 g water contain 40 g KCl

1 g water contain = (40/180) g KCl

100 g water contain = (40×100/180) g

                   = 2.22 g KCl

                   

Hence the concentration in mass percentage of KCl = 2.22 %